# High-dimensional stochastic finite volumes using the tensor train format

### v1

@@ 4.2 Computing expectation and variance
Once the solution in the mixed format $\bar{U}$ is known, we compute its expectation and variance. For simplicity, we describe the computation for two stochastic variables $\xi_{1},\xi_{2}$. The expectation and variance of the solution $u$ at time $t$ and point $x$ are
$\mathbb{E}[u(t,x)]$ $=\int_{\Omega_{\bm{\xi}}}u(t,x,\xi_{1},\xi_{2})p_{\bm{\xi}}({\bm{\xi}})\;\text{d}{\bm{\xi}},$ (32)
@@
This operation can be seen as a contraction of the tensor $\bar{u}_{i,j_{1},j_{2}}^{n}$ with the tensor $p(\xi_{j_{1}})p(\xi_{j_{2}})$. Moreover, the tensor $p(\xi_{j_{1}})p(\xi_{j_{2}})$ is a tensor train with rank 1 and cores $p(\xi_{j_{1}})$ and $p(\xi_{j_{2}})$. With the mixed-format approximation $\bar{U}_{i}^{n}$, the tensor contraction can be computed in the tensor train format. For the variance, we first compute the squared solution in the mixed format by operation on tensor trains, then we compute its expectation as above and finally obtain the variance from Equation (32).
@@ 5 Numerical results
In this section, we apply the method to two hyperbolic problems: the Riemann problem associated to Burgers’ equation, and the Sod problem for the Euler equations. The first problem is used to investigate some properties of the method. We study two performance indicators: the convergence order in the expectation and variance, and the number of tensor entries. We also evaluate the influence of some parameters on the performance of the algorithm, namely the TT-rank and the tolerance for the cross approximation algorithm. Finally, we show that the method can scale up for a larger number of stochastic dimensions. The second problem showcases the method’s applicability to a system of equations, and its ability to capture effects such as shock, rarefaction, and contact waves.
@@ 5.1 Stochastic Burgers’ equation
We consider the Burgers’ equation with uncertain initial conditions,
@@
Let ${\bf v}_{L}=(v_{L,1},\dots,v_{L,m})^{\top}$ and ${\bf v}_{R}=(v_{R,1},\dots,v_{R,m})^{\top}$. For now, the number $m$ of random variables is arbitrary. The value of $m$ and of the coefficients ${\bf v}_{L},{\bf v}_{R}$ will be given in each subsection. We focus on the case where the solution is a shock wave, hence, we assume that for all realizations of ${\bm{\xi}}=(\xi_{1},\dots,\xi_{m})$ it holds
@@
5.1.1 Convergence study and influence of algorithmic parameters
The random variables are assumed independent and $\xi_{\ell}\sim\mathcal{U}(0,1)$, for all $\ell\in\{1,\dots,m\}$. We fix $m=3$, $(v_{L,1},v_{L,2},v_{L,3})=(0.1,0,-0.1)$, $(v_{R,1},v_{R,2},v_{R,3})=(0.1,-0.1,0)$. The initial condition then reads
@@ 5.1.1 Convergence study and influence of algorithmic parameters
We investigate the influence of some parameters on the performance of the algorithm. Various performance indicators can be used, here we focus on three values: the $L^{1}$-error between the approximated and exact solution expectation, the $L^{1}$-error between the approximated and exact solution variance, and the compression ratio for the solution at the final time. The compression ratio is defined as the number of entries for the mixed format divided by the number of entries for the full tensor. The relative $L^{1}$-error in expectation at final time $t^{n}$ is defined by
$|\mathbb{E}|_{L^{1},\text{rel}}=\frac{|\mathbb{E}_{\text{TT}}-\mathbb{E}_{\text{ex}}|_{L^{1}}}{|\mathbb{E}_{\text{ex}}|_{L^{1}}},\quad|\mathbb{E}|_{L^{1}}=\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}\mathbb{E}[\bar{u}_{i}^{n}].$ (39)
Here the expectations for the exact and approximated solution are computed with
$|\mathbb{E}_{TT}-\mathbb{E}_{ex}|_{L^{1}}=\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}\left(\mathbb{E}[U_{i}^{n}]-\int_{x_{i-1/2}}^{x_{i+1/2}}\mathbb{E}[u(x,t)]\right).$ (40)
@@
For the integration of the exact expectation along $x$, we use the midpoint rule. The expectation for the tensor train $\mathbb{E}[U_{i}^{n}]$ is computed as described in Section 4.2. The expectation and the variance of the exact solution are given in Appendix B. Since their expression requires nontrivial integration, the integrals are numerically approximated using the scipy.integrate package [19].
@@
The investigated parameters are the number of stochastic and spatial cells, the TT-rank, and the tolerance for the cross approximation function. We give here some explanation for each parameter. As before, the number of cells in the spatial direction is denoted by $N_{x}$. Each stochastic direction has the same discretization, and we denote by $N_{\xi}$ the number of cells for each stochastic direction. For example, in the case $N_{x}=N_{\xi}=100$, the spatial-stochastic grid is made of $100^{4}$ cells. The TT-rank $r$ corresponds to the maximal rank allowed for each rounding operation. The tolerance $\varepsilon$ is a parameter related to the cross approximation algorithm. The iterative algorithm terminates once the residual on the selected entries is less than $\varepsilon$. It should be noted that this tolerance does not give a global error on the tensor approximation, so its influence on the total error of the algorithm is not known.
Figure 1 shows the case $N_{x}=N_{\xi}\in\{40,\dots,100\}$, and Figure 2 shows the case $N_{\xi}=20$ and $N_{x}\in\{40,\dots 100\}$. In addition to the results shown in these figures, we also tried other values for the TT-rank and the tolerance. For the TT-rank, we tried with $r=10,r=100$ and setting no maximal ranks, but this led to an excess of memory for high $N_{x}$. For the tolerance, we tried with $\varepsilon=10^{-6}$, as it is the default value in the library tntorch. Then the algorithm became much slower, and the cross approximation failed to converge in many cases. As we see in the results, it is not clear whether a smaller tolerance improves the accuracy of the solution.
Figure 1: $L^{1}$-error of the expectation and variance, and number of parameters for a shock wave for $N_{x}=N_{\xi}$. The curves for the same $r$ and different $\varepsilon$ are superposed.
Figure 2: $L^{1}$-error of the expectation and variance, and a number of parameters for a shock wave for $N_{\xi}=20$. The curves for the same $r$ and different $\varepsilon$ are superposed.
From Figures 1 and 2 we make the following observations:
  • The $L^{1}$-errors are higher for odd $N_{x}$ than for even $N_{x}$. This might be caused by the location of shock relative to the cell boundary. It should also be noted that for small $N_{x}$, i.e. $N_{x}<60$, similar oscillations can be seen in the exact solution.
  • The convergence order is approximately 1. For a non-stochastic problem, a MUSCL scheme with a midpoint integration and a central Euler is expected to converge of second order. However, to our knowledge, no convergence order has been proved for the solution expectation of a stochastic problem. It would also be interesting to investigate how tensor approximation alters the convergence order.
  • Changing the tolerance $\varepsilon$ does not seem to have a noticeable effect.
  • A larger TT-rank $r$ yields a lower $L^{1}$-error for the variance. The $L^{1}$-error for the expectations is not strictly smaller, but its oscillation in $N_{x}$ is smaller.
  • The compression ratio is globally very small, even for $r=5$.
  • For a fixed $N_{x}$, the version with higher $N_{\xi}$ is slightly more accurate. Increasing $N_{\xi}$ does not seem to have an impact on the convergence order in $x$. We can assume that here, the loss of convergence order in $x$ is not related to the discretization in $N_{\xi}$.
@@ 5.1.2 Scaling study
The random variables are assumed independent and $\xi_{\ell}\sim\mathcal{U}(0,1)$, for all $\ell\in\{1,\dots,m\}$. We fix $N_{x}=N_{\xi}=20,r=5,\varepsilon=0.1$ and vary the number of stochastic dimensions $m$. Here the coefficients ${\bf v}_{R},{\bf v}_{L}$ are $v_{R,\ell}=0.1,~{}v_{L,\ell}=-0.1$ for $\ell\in\{1,\dots,m\}$. This ensures that the solution is a shock wave, no matter the number of stochastic dimensions. Thus, the expectation will exhibit a discontinuity at around $x=0$, see Figure 4.
Figure 3: Number of entries for the solution at the final time, compared to the number of entries for the full tensor.
Figure 3 shows the number of entries for the solution at the final time and compare it with the number of entries for the full tensor. The number of entries for the mixed format is polynomial with respect to the number of stochastic dimensions, whereas it is exponential for the full format. When running the experiment, we noticed that the cross approximation algorithm did not always converge. The cross approximation failed to converge more frequently with increasing dimensions. To check how much the convergence of the cross approximation alters the solution, we show in Figure 4 the expectation of the solution. The numerical solution seems not to be affected by the failed convergence: a shock is still visible at the expected location.
Figure 4: Expectation of the solution as a function of $x$, for various number of stochastic dimensions $m$.
5.2 Euler equations
The last experiment is the Sod problem for the Euler equations. We use this example to show that the algorithm can be extended to systems, and that the approximated solution can capture features such as shocks and contact waves typical for nonlinear hyperbolic problems. We also investigate the influence of the mesh size and the rank on the solution.
@@ 5.2.1 Problem description
Here $\rho$ is the density, $u$ the velocity, $E$ the total energy, $p$ the pressure. For an ideal fluid with a specific heat ration $\gamma=1.4$, the pressure is given by
$p=(\gamma-1)(\rho E-\frac{\rho u^{2}}{2}).$ (42)
Note that compared to the Burgers’ equation, the flux of the Euler equations consists in non-polynomial functions. Hence, cross approximation is required for the flux evaluation. We consider the Sod shock tube problem for $x\in[0,1]$ with free boundary conditions and with $m=3$ stochastic directions,
@@ 5.2.2 Parameter study
The Sod problem is used to investigate the influence of some parameters on the solution. We focus here on the mesh size in the spatial and stochastic directions, and on the maximal rank. The tolerance for cross approximation, that was varied in the Burgers case, is fixed here at $\varepsilon=0.1$. Since there is no analytical solution known for the Sod problem, we focus here on the qualitative results. We plot the expectation and variance of the primitive variables (density, velocity and pressure) at final time $t=0.2$. For comparison, the expectation and variance are also computed using a Monte-Carlo method with the library MultiWave [20]: an adaptive discontinuous Galerkin solver was used with $L=6$ refinement levels corresponding to a uniformly refined grid with 192 cells and polynomial elements of degree 2, i.e., third-order scheme; the expectation and variance were computed with 125,000 samples. Details can be found in [21, 16, 17] where similar Monte Carlo Simulations were performed.
@@
Figures 5 and 6 respectively show the expectation and variance of the primitive variable for $r=5,N_{\xi}=20$ and for various mesh refinement in the physical space $N_{x}$. For the expectation, there are little differences between $N_{x}=80$ and $N_{x}=160$, so the scheme seems to converge. Moreover, the general shape is in globally in good agreement with the MC results. However, in transition between the contact wave and the shock wave at $x\in[0.65,0.85]$, there are non-physical oscillations, which will be further investigated in the next paragraph. We also note that the results with MC are sharper, which is probably due to the fact that a third-order DG solver was used. For the variance, in some areas convergence is slower and a local refinement could be beneficial.
Figure 5: Expectation of $\rho,u,p$ for $r=5,N_{\xi}=20$ and $N_{x}\in\{20,40,80,160\}$. The expectation obtained with a Monte-Carlo (’MC’) is in dotted line.
Figure 6: Variance of $\rho,u,p$ for $r=5,N_{\xi}=20$ and $N_{x}\in\{20,40,80,160\}$. The variance obtained with a Monte-Carlo (’MC’) is in dotted line.
Figure 7 shows the expectation of the primitive variables for $N_{x}=160,N_{\xi}=20$ and for various values of the rank $r$. All the tested ranks yield mostly the same result, which supports the idea that the solution can be represented in most cells with low-rank tensors. Only the zone $x\in[0.65,0.85]$, between the contact wave and the shock, shows differences for various ranks, see Figure 8. Increasing the rank is here necessary to reduce the oscillations. Here, we see that increasing locally the rank would improve the solution while keeping computational cost low. We also tried ranks lower than 5, but this yielded non-physical solutions.
Figure 7: Expectation of $\rho,u,p$ for $N_{x}=160,N_{\xi}=20$, and $r\in\{5,15\}$
Figure 8: Zoom on expectation of $u$ for $N_{x}=160,N_{\xi}=20$, and $r\in\{5,15\}$
@@
Figure 9 and 10 respectively show the expectation and variance of the primitive variable for $N_{x}=160,r=5$ and for various mesh refinement in the stochastic space $N_{\xi}$. Here, the number of stochastic cells does not seem to have a significant influence, which is probably due to the choice of a uniform probability distributions for each random variable.
Figure 9: Expectation of $\rho,u,p$ for $N_{x}=160,r=5$ and $N_{\xi}\in\{20,40,80\}$.
Figure 10: Variance of $\rho,u,p$ for $N_{x}=160,r=5$ and $N_{\xi}\in\{20,40,80\}$.

### v2

@@ 4.2 Computing expectation and variance
Once the solution in the hybrid format $\bar{U}$ is known, we compute its expectation and variance. For simplicity, we describe the computation for two stochastic variables ${\bm{\xi}}=(\xi_{1},\xi_{2})$. The expectation and variance of the solution $u$ at time $t$ and point $x$ are
$\mathbb{E}[u(t,x)]$ $=\int_{\Omega_{\bm{\xi}}}u(t,x,\xi_{1},\xi_{2})p_{\xi_{1},\xi_{2}}(\xi_{1},\xi_{2})\;\text{d}\xi_{1}\;\text{d}\xi_{2},$ (33)
@@
This operation can be seen as a contraction of the tensor $\bar{u}_{i,j_{1},j_{2}}^{n}$ with the tensor $p(\xi_{j_{1}})p(\xi_{j_{2}})$. Moreover, the tensor $p(\xi_{j_{1}})p(\xi_{j_{2}})$ is a tensor train with rank 1 and cores $p(\xi_{j_{1}})$ and $p(\xi_{j_{2}})$. With the solution in the hybrid format $\bar{U}_{i}^{n}$, the tensor contraction can be computed in the tensor train format. For the variance, the squared solution is first computed in the hybrid format by operation on tensor trains, then its expectation is obtained as above. The variance is then computed using Equation (34).
4.3 The fullly-TT format for the MUSCL scheme
We compare the proposed approach with the full-TT approach. For this, the method introduced in [7] is stated for the MUSCL scheme. The core idea is to rewrite the scheme (15)-(20) in terms of global matrices. Then, every computational operation is applied to the full TT approximation of the solution, using either TT algebraic operations or cross-approximation.
We illustrate the idea by describing the reconstruction step with global matrices. For simplicity, the operation is first described in the scalar case and a single spatial variable and without any stochastic variables. Let $\bar{u}^{n}=(u_{1}^{n},\dots u_{N_{x}}^{n})$ be the solution vector. We introduce the identity matrix ${\bf I}\in\mathbb{R}^{N_{x}\times N_{x}}$ and the matrices ${\bf L},{\bf R}\in\mathbb{R}^{N_{x}\times N_{x}}$ defined by
${\bf L}=\begin{pmatrix}0&0&0&\cdots&0\\ 1&0&0&\cdots&0\\ 0&1&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&0\\ 0&0&\cdots&1&0\\ \end{pmatrix}\text{ and }{\bf R}=\begin{pmatrix}0&1&0&\cdots&0\\ 0&0&1&\cdots&0\\ 0&0&0&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&0&1\\ 0&0&\cdots&0&0\\ \end{pmatrix}$.
The vector of approximate derivatives defined in (18) is obtained as matrix vector multiplication with the defined matrices, i.e.,
$(u_{x})^{n}=\frac{1}{\Delta x}\text{minmod}\left(({\bf I}-{\bf L})\bar{u}^{n},({\bf R}-{\bf I})\bar{u}^{n}\right),$ (36)
and the MUSCL scheme then computes the right and left states at the cell interfaces, denoted by $u^{n\pm}=(u_{3/2}^{n\pm},\dots,u_{N_{x}+1/2}^{n\pm})$ and defined with (16),
$u^{n\pm}=\bar{u}^{n}\pm\frac{\Delta x}{2}(u_{x})^{n}.$ (37)
In the case of $m$ stochastic variables, the solution is approximated as a tensor train using the representation (25). Now, the previous computations are repeated on the tensor train format. Since the operation of shifting an index to the right or to the left (given by the previously defined matrices) is only applied to the spatial indices, only the spatial core $G_{0}$ of the representation is affected. For example, the right-shifted core $\tilde{G}_{0}$ is given by $\tilde{G}_{0}={\bf R}G_{0}$, and the tensor-train for the right-shifted solution is
$\bar{U}^{n}(i+1,{\bm{j}})=\sum_{\alpha_{1}\dots\alpha_{m}=1}^{r_{1}\dots r_{m}}\tilde{G}_{0}(i,\alpha_{1})G_{1}(\alpha_{0},j_{1},\alpha_{1})G_{2}(\alpha_{1},j_{2},\alpha_{2})\cdots G_{m}(\alpha_{m},j_{m}).$ (38)
The approximate derivative in the TT format is thus obtained by computing the left- and right-shifted cores and by applying the minmod limiter to the resulting tensor trains. The later requires e.g. cross–approximation on the tensor train format. The full-TT approximation in the MUSCL scheme is then summarized in Algorithm 3. The remaining operations are linear on the tensor train and their description is omitted. Finally, expectation and variance are obtained by integration that is done by contracting the solution tensor with the rank-one tensor of integration weights.
@@ 5 Numerical results
In this section, we apply the method to three hyperbolic problems: the Riemann problem associated to Burgers’ equation, the Sod problem and the Shu-Osher problem for the Euler equations. For the first problem, we investigate the influence of some parameters on the algorithm accuracy, and compare it to the full-TT format. We also show that the method can scale up for a larger number of stochastic dimensions. The second problem showcases the method’s applicability to a system of equations. A parameter study and a comparison with the full-TT format are presented. Finally, the last test case shows the method’s ability to handle a complex interaction problem.
@@ 5.1 Stochastic Burgers’ equation
We consider the Burgers’ equation with uncertain initial conditions
@@
The random variables are assumed independent with $\xi_{\ell}\sim\mathcal{U}(0,1)$, for all $\ell\in\{1,\dots,m\}$. Computations are performed for $\Omega_{x}=(-1,1)$ until final time $T=0.35$, and we use the CFL number $\alpha=0.45$. Let ${\bf v}_{L}=(v_{L,1},\dots,v_{L,m})^{\top}$ and ${\bf v}_{R}=(v_{R,1},\dots,v_{R,m})^{\top}$. For now, the number $m$ of random variables is arbitrary. The value of $m$ and of the coefficients ${\bf v}_{L},{\bf v}_{R}$ will be given in each subsection. We focus on the case where the solution is a shock wave, hence, we assume that for all realizations of ${\bm{\xi}}=(\xi_{1},\dots,\xi_{m})$ it holds
@@
5.1.1 Convergence study and influence of parameters
We choose $m=3$ and take the following initial conditions
@@ 5.1.1 Convergence study and influence of parameters
We investigate the influence of some parameters on the performance of the algorithm.
The number of stochastic and spatial cells, the TT-rank, and the tolerance for TT approximations are varied, and the relative $L^{1}$-error of the expectation, the relative $L^{1}$-error of the variance, and the number of coefficients for the hybrid format at final time are computed. The relative $L^{1}$-error for the expectation at final time $T$ is defined by
$|\mathbb{E}|_{L^{1},\text{rel}}=\frac{|\mathbb{E}_{\text{TT}}-\mathbb{E}_{\text{ex}}|_{L^{1}}}{|\mathbb{E}_{\text{ex}}|_{L^{1}}}.$ (43)
The exact expectation along $x$ is approximated using the midpoint rule, and the relative error is computed with
$|\mathbb{E}|_{L^{1}}=\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}\mathbb{E}[\bar{u}_{i}^{n}],\qquad|\mathbb{E}_{TT}-\mathbb{E}_{\text{ex}}|_{L^{1}}=\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}\left(\mathbb{E}[U_{i}^{n}]-\int_{x_{i-1/2}}^{x_{i+1/2}}\mathbb{E}[u(x,t)]\right).$ (44)
@@
The relative $L^{1}$-error for the variance $|\text{var}|_{L^{1},\text{rel}}$ is defined similarly. The expectation for the tensor train $\mathbb{E}[U_{i}^{n}]$ is computed as described in Section 4.2. The expectation and the variance of the exact solution are given in Appendix B. Since their expression requires nontrivial integration, the integrals are numerically approximated using the scipy.integrate package [19].
@@
The number of cells in the spatial direction is denoted by $N_{x}$. Each stochastic direction has the same discretization, and we denote by $N_{\xi}$ the number of cells for each stochastic direction. For example, in the case $N_{x}=N_{\xi}=160$, the spatial-stochastic grid is made of $160^{4}$ cells. The TT-rank $r_{\text{max}}$} corresponds to the maximal rank allowed for each rounding operation. The tolerance $\varepsilon$ is a parameter related to the cross approximation algorithm and the rounding operations. It should be noted that this tolerance does not give a global error on the tensor approximation, so its influence on the total error of the algorithm is not known. Table 1 shows the tested parameters. The maximal rank actually observed for each simulation is denoted by $r$ in the table.
We first note that for the expectation, except for the case $\varepsilon=0.1$, all combinations of $\varepsilon$ and $r_{\text{max}}$} result in quite similar convergence curves. When $\varepsilon<0.1$, the convergence order is close to one for the expectation and one half for the variance. Those orders are in agreement with the one-half order predicted by the theory [8]. For the case $\varepsilon=0.1$, the TT approximation is not precise enough, and the expected convergence rate is not reached. This behavior was also observed in [7] for the full-TT discretization. The lowest rank ${r_{\text{max}}}=1$ seems to be as good as larger ranks for the expectation, but not for the variance. From Table 1, we see that the maximal rank $r_{\text{max}}$} is more often reached the larger $N_{x}$ and the smaller $\varepsilon$. Finally, we note that during the simulations, warnings indicated that the cross-interpolation was not converging, even when $r_{\text{max}}$} is not reached and even though the final errors are reasonable.
The convergence study was also done by varying only $N_{x}$ and fixing $N_{\xi}=20$. Figure 2 shows the resulting convergence curves. They are very similar to the case $N_{\xi}=N_{x}$, indicating that the convergence rate is mostly driven by the accuracy in $x$.
$\varepsilon=10^{-1}$ $\varepsilon=10^{-3}$ $\varepsilon=10^{-5}$
$N_{x}$ $r_{\text{max}}$} $r$ (h) $r$ (f) $N_{x}$ $_{\text{max}} rr$ (h) $r$ (f) $N_{x}$ $_{\text{max}} rr$ (h) $r$ (f)
1 1 1 1 1 1 1 1 1
20 5 2 1 20 5 2 3 20 5 5 5
10 2 1 10 2 3 10 8 10
30 2 1 30 2 3 30 12 13
1 1 1 1 1 1 1 1 1
40 5 2 1 40 5 5 4 40 5 5 5
10 2 1 10 6 4 10 10 10
30 2 1 30 8 5 30 30 30
1 1 1 1 1 1 1 1 1
80 5 2 1 80 5 5 5 80 5 5 5
10 2 1 10 9 6 10 10 10
30 2 1 30 12 7 30 30 30
1 1 1 1 1 1 1 1 1
160 5 2 1 160 5 5 5 160 5 5 5
10 2 1 10 10 7 10 10 10
30 2 1 30 15 7 30 30 30
Table 1: Parameters tested for Sections 5.1.1. The maximal rank observed for the hybrid and the full-TT format in the case $N_{\xi}=N_{x}$ are respectively denoted by $r$ (h) and $r$ (f).
Figure 1: Relative $L^{1}$-errors of the expectation and variance for the shock wave, computed with $N_{\xi}=N_{x}$ using the hybrid method.
Figure 2: Relative $L^{1}$-errors of the expectation and variance for the shock wave, computed with $N_{\xi}=20$ using the hybrid method.
We also compare the performance of the hybrid format to the full-TT format. Figure 3 shows the convergence curve of the hybrid and the full-TT format using the values of $\varepsilon$ and $r_{\text{max}}$} given in Table 1. The case $\varepsilon=0.1$ yields better results for the hybrid format than for the full-TT format. For smaller $\varepsilon$, both methods yield comparable results. In this case, the full-TT method is more efficient since it requires fewer tensor coefficients: for the same maximal rank $r$, the hybrid method requires $N_{\xi}$ times more parameters. Figure 4 shows how the number of coefficients at final time increases with $N_{x}$ for each method, compared to the theoretical number of parameter for the full discretization. We use $\varepsilon=10^{-3},{r_{\text{max}}}=5$ and $N_{\xi}=N_{x}$. As expected, the full-TT format has a lower number of coefficients than the hybrid format, and both methods provide a significant compression compared to the full tensor.
Figure 3: Relative $L^{1}$-error of the expectation for the shock wave, using the full-TT (plain line) and hybrid (dashed line) formats.
Figure 4: Number of coefficients of the solution at the final time: hybrid format, full-TT format, and theoretical value for the full tensor.
@@ 5.1.2 Scaling study
We now investigate the scalability of the method. We fix $N_{x}=40,N_{\xi}=20,{r_{\text{max}}}=5,\varepsilon=0.01$ and vary the number of stochastic dimensions $m$. The coefficients ${\bf v}_{R},{\bf v}_{L}$ are $v_{R,\ell}=0.1,~v_{L,\ell}=-0.1$ for $\ell\in\{1,\dots,m\}$. This ensures that the solution is a shock wave, no matter the number of stochastic dimensions. Thus, the expectation will exhibit a discontinuity at around $x=0$, see Figure 6.
Figure 5: Number of coefficients of the solution at the final time for increasing stochastic dimensions: hybrid format and theoretical value for the full tensor
Figure 5 shows the number of entries for the solution at the final time and compare it with the number of entries for the full tensor. The number of entries for the hybrid format is polynomial with respect to the number of stochastic dimensions, whereas it is exponential for the full format. When running the experiment, we noticed that the cross approximation algorithm did not always converge. The cross approximation failed to converge more frequently with increasing dimensions. To check how much the convergence of the cross approximation alters the solution, we show in Figure 6 the expectation of the solution. The numerical solution seems not to be affected by the failed convergence: a shock is still visible at the expected location.
Figure 6: Expectation of the solution as a function of $x$, for various numbers of stochastic dimensions $m$.
5.2 Sod problem
The second experiment is the Sod problem for the Euler equations. We use this example to show that the algorithm can be extended to systems, and that the approximated solution can capture features such as shocks and contact waves typical for nonlinear hyperbolic problems. We also investigate the influence of the mesh size and the rank on the solution.
@@ 5.2.1 Problem description
Here $\rho$ is the density, $u$ the velocity, $E$ the total energy, $p$ the pressure. For a perfect gas with specific heat ratio $\gamma=1.4$, the pressure is given by
$p=(\gamma-1)\left(\rho E-\frac{\rho u^{2}}{2}\right).$ (46)
Note that compared to the Burgers’ equation, the flux of the Euler equations consists of non-polynomial functions. Hence, cross approximation is required for the flux evaluation. We consider the Sod shock tube problem for $x\in(0,1)$ with free boundary conditions and with $m=3$ stochastic directions,
@@ 5.2.2 Parameter study
We investigate the influence of $N_{x}$, $N_{\xi}$ and $r_{\text{max}}$} on the solution. The tolerance for TT approximations, that was varied in the Burgers case, is fixed here at $\varepsilon=0.01$. Since there is no analytical solution known for the stochastic Sod problem, we focus here on the qualitative results. We plot the expectation and variance of the primitive variables at final time $T=0.2$. For comparison, the expectation and variance are also computed using a Monte-Carlo (MC) method with the library MultiWave [20]: an adaptive discontinuous Galerkin solver was used with $L=6$ refinement levels corresponding to a uniformly refined grid with 192 cells and polynomial elements of degree 2, i.e., third-order scheme; the expectation and variance were computed with 125,000 samples. Details can be found in [21, 16, 17] where similar Monte-Carlo simulations were performed.
@@
Figures 7 and 8 respectively show the expectation and variance of the primitive variable for $r=5,N_{\xi}=20$ and for various mesh refinement in the physical space $N_{x}$. For the expectation, there are little differences between $N_{x}=80$ and $N_{x}=160$ indicating convergence of the scheme. The general shape is globally in good agreement with the MC results. In transition between the contact wave and the shock wave at $x\in[0.65,0.85]$, there are non-physical oscillations. These oscillations are also present in the deterministic Euler equations when solving with a MUSCL scheme and a leap-frog time-stepping. Hence, they are probably not related to the stochastic problem, and will not be further investigated in this work. The MC results are sharper, as they are obtained with a third-order DG solver.
Figure 7: Expectation of $\rho,u,p$ for ${r_{\text{max}}}=5,N_{\xi}=20$ and $N_{x}\in\{20,40,80,160\}$. The MC result (dashed line) is plotted for reference.
Figure 8: Variance of $\rho,u,p$ for ${r_{\text{max}}}=5,N_{\xi}=20$ and $N_{x}\in\{20,40,80,160\}$. The MC result (dashed line) is plotted for reference
To study the influence of the rank, we set $N_{x}=160,N_{\xi}=20$ and vary the maximal rank $r_{\text{max}}$}. The expectation and variance of $\rho,u,p$ are shown in Figure 9 and in Figure 10 respectively. For $r_{\text{max}}$} larger than 5, the results are very similar. We also tested ranks lower than 5, but the solution became non-physical after some time. Figure 11 shows the actual maximal rank at final time in each spatial cell for the case $N_{x}=160,N_{\xi}=20$ and ${r_{\text{max}}}=10$. In most cells, the ranks are very low – less than three – and the large ranks concentrate in few cells around the shock. Hence, large ranks are needed only in localized regions. Moreover, they can provide information on the spatial structure of the solution.
Figure 9: Expectation of $\rho,u,p$ for $N_{x}=160,N_{\xi}=20$, and ${r_{\text{max}}}\in\{5,10,15,20\}$.
Figure 10: Variance of $\rho,u,p$ for $N_{x}=160,N_{\xi}=20$, and ${r_{\text{max}}}\in\{5,10,15,20\}$.
Figure 11: Maximal rank observed at final time for each spatial cell in the hybrid format of $\rho,u,p$ for $N_{x}=160,N_{\xi}=20,{r_{\text{max}}}\in\{5,10,15,20\}$.
@@
Finally, we fix $N_{x}=160,r=5$ and vary $N_{\xi}$. Figure 12 shows the expectation of $\rho,u,p$. The number of stochastic cells does not seem to have a significant influence, which is probably due to the choice of a uniform probability distribution for each random variable. It should be noted that the case $N_{\xi}=40$ yields non-physical solutions, hence it is not included here.
Figure 12: Expectation of $\rho,u,p$ for $N_{x}=160,{r_{\text{max}}}=5$ and $N_{\xi}\in\{20,80\}$.
5.2.3 Comparison with the full-TT format
The results obtained with the hybrid format are compared with the full-TT formats. Figure 13 shows the expectation of $\rho,u,p$ using $N_{x}=80,N_{\xi}=20,\varepsilon=10^{-3},{r_{\text{max}}}=10$ for both formats. The MC result is used as reference. We see that the hybrid format is slightly more accurate than the full-TT format, in particular for the density and around the shock. However, with our current implementation, the hybrid format ran for much longer than the full-TT format: it took approximately two hours to complete, against ten minutes for the full-TT format. The number of coefficients at final time is also much larger for the hybrid format. Those two aspects could be greatly improved by improving the finite volume scheme, using e.g. a higher-order method and a non-uniform mesh in the space direction.
Figure 13: Expectation of $\rho,u,p$ for the full-TT format and the hybrid format for $N_{x}=80,N_{\xi}=20$ and ${r_{\text{max}}}=10$. The MC result (dashed line) is plotted for reference.